OpenStudy (anonymous):

Write 32x^5y^5 with only one exponent. Use parentheses.

OpenStudy (anonymous):

(2xy)^5

OpenStudy (anonymous):

32 = 2 x 2 x 2 x 2 x 2 = 2^5

OpenStudy (anonymous):

Remember in indices, a^m x b^m = (ab)^m

OpenStudy (anonymous):

5y^5 is really 5yx5yx5yx5yx5y, which is 5^5 y^5, which is 3125y^5